OpenStudy (anonymous):

Rosanne drops a ball from a height of 400ft. Find the ball's average height nd its average velocity between the ime it is dropped and the time it strikes the ground.

OpenStudy (dumbcow):

From the acceleration of gravity being -32 ft/s^2 and initial velocity is zero we know that: h(t) = -16t^2 + 400 v(t) = -32t To find time it hits ground, solve h(t) = 0 16t^2 = 400 --> t = 5 For average height and avg velocity, use integrals \[\frac{1}{5}\int\limits_{0}^{5} h(t) dt\] \[\frac{1}{5}\int\limits_{0}^{5} v(t) dt\]

OpenStudy (anonymous):

Thank you very much.